OpenStudy (anonymous):

solve for X (x-1)(x-2)(x-3)-0

OpenStudy (anonymous):

(x-1)(x-2)(x-3)=0 it will = 0 only if atleast any one term=0 x-1=0 or x-2 =0 or x-3 =0 x=1 or x=2 or x=3